OpenStudy (anonymous):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (anonymous):

@gibbs

OpenStudy (gibbs):

A

OpenStudy (anonymous):

can you help me with more?

OpenStudy (gibbs):

-1 + 1=2 like sighs add and use that sighn unlike subtract and use the sighn of the bigger

OpenStudy (anonymous):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

OpenStudy (gibbs):

C

OpenStudy (anonymous):

but there are multiple answers

OpenStudy (gibbs):

C and D

OpenStudy (gibbs):

the first one is A and B

OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (anonymous):

oh whoops one sec

OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (anonymous):

@gibbs

OpenStudy (gibbs):

um they are all corect exept D

OpenStudy (anonymous):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2}

OpenStudy (gibbs):

C D

OpenStudy (anonymous):

Numbers p and q are positive integers. Which statements are always true? Choose all answers that are correct. A. p + q is a positive integer B. p – q is a negative integer C. p • q is a positive integer D. is a positive integer

OpenStudy (gibbs):

A B C

OpenStudy (anonymous):

Is the set of multiples of 5 closed under division? Explain why or provide a counterexample if not. A. Yes, because the quotient of any two multiples of 5 is also a multiple of 5. B. Yes, because the quotient of any two multiples of 5 is also a multiple of 10. C. No, and a counterexample is 100 ÷ 5 = 20. D. No, and a counterexample is 25 ÷ 50 = .

OpenStudy (gibbs):

B

OpenStudy (anonymous):

@gibbs dude i got 20%

OpenStudy (gibbs):

How i googled most of the answeres

OpenStudy (anonymous):

. . .